Please help me solve this. Its not homework, just practice for my test.
82+X/4=165

Answer by checkley71(8403)

82+x/4=165
x/4=165-82
x/4=83 now cross multiply
x=4*83
x=332 answer.
PROOF:
82+332/4=165
82+83=165
165=165